what is the answer to:
*what is the slope of a line with points (9,2) & (3,10)

Answer by mananth(16946)

*what is the slope of a line with points (9,2) & (3,10)

slope m=Y2-y1/x2-x1
..
m= 10-2 / 3-9
m=8/-6
m=-4/3